Air Shangrila sells to both tourist and business travellers on its single route. Tourists always stay over on Saturday nights, while business travellers never do. The weekly demand function of tourists is QdT = 6000-10P, and the weekly demand function of business travellers is QdB=1000-P. If the marginal cost of a ticket is $200, what prices should Air Shangrila set for its tourist ticket and its business ticket? What would be the elasticity of demand for the two groups at that price?
Answer #1

Case-I- Tourist Qdt = 6000-10P P= Gooo-Q 10 - & TR = /6000-Q) Q (6000-2) Total Revenue = P.Q IMR 6000 - 20 10 MC= $200 degreeCase -II - Business tickets Qde = 1000-P x TP = 1000-Q Q) * 2 Revenne = 1000-2Q Marginal Tми IMR = 1000-2Q MC= $200 be The co